If a particle os moving along yaxis the position varied with time t as y=t^2-3t+1 where y is in meter and t is in second the distance traveled by particle on first 2 second`s

Grade:12th pass

1 Answers

lj
6 years ago
We have to check whether velocity of particle becomes zero in its path in time interval 2 sec. And whether its velocity`s direction changes So, dy/dt = 2t-3=vv=0 when t=3/2Put t=3/2,y=-7/2Put y=2,y=-1This, distance = 7/2+(7/2-1)=6